OpenStudy (anonymous):

The volume of a cone is cubic kilometers. If the radius of the base is 5 kilometers, what is the height of the cone? Use for . Show your work and round to the nearest tenth.

hero (hero):

Curry, it's all yours

OpenStudy (curry):

aight

OpenStudy (anonymous):

plz help me

OpenStudy (curry):

u have to provide the volume to find the height

OpenStudy (curry):

or else it would be V/(25pi/3) = h

OpenStudy (anonymous):

it says use 3.14 for pi

hero (hero):

It doesn't say what the volume is?

OpenStudy (curry):

nope

OpenStudy (anonymous):

the volume is 300 cubic kilometers

OpenStudy (curry):

V/ 26.17 = h

hero (hero):

There we go :)

OpenStudy (curry):

11.46 = h

hero (hero):

300/26.17 =h

OpenStudy (curry):

hero check my work please

OpenStudy (anonymous):

thank you sooo much (:

OpenStudy (curry):

np